OpenStudy (anonymous):

An express train traveled 600 miles in the same amount of time for a freight train to travel 360 miles. The rate of the express train was 20 miles faster than the rate of the freight train. Find the rate of each train.An express train traveled 600 miles in the same amount of time for a freight train to travel 360 miles. The rate of the express train was 20 miles faster than the rate of the freight train. I solved all the way to 60r=360r+7200

OpenStudy (anonymous):

600/(r+20)=360/r then 600r=360r+7200 then r =30 the rates are 30 and 50